### v1

# Resolving the Exploitation-Exploration Dilemma in Evolutionary Algorithms: A Novel Human-Centered Framework

Abstract
Evolutionary Algorithms (EAs) are powerful tools for tackling complex computational problems, yet effectively managing the exploitation and exploration dynamics—crucial for robust search navigation—remains a persistent challenge for EA designers, and leads to the so-called exploration-exploitation dilemma. In this paper, a new human-centered framework is proposed to resolve this dilemma. Unlike the traditional approach, the search process will not be compromised of a single-phase nor the decision-maker tuning efforts will be distributed among the algorithm’s traditional parameters such as defining new evolutionary operators internal to the algorithm to influence its search navigation. Instead, a human-centered two-phase search process, compromised of a global search phase followed by a local phase will be utilized. In this framework, the designer plays the central role in directing the algorithm’s search navigation through the focused tuning efforts of a new Search Space Size Control parameter external to the algorithm which proves itself to be the dominant parameter in-effect to the algorithm’s effective search navigation.
The framework is applicable to any search algorithm. We demonstrate its effectiveness on 14 well-known benchmark problems in unconstrained optimization.
Keywords Exploration-Exploitation Dilemma  $\cdot$ Evolutionary Algorithms  $\cdot$ Search Space Size Control Parameter  $\cdot$ Human-Centered Two-Phase Search Framework  $\cdot$ Global Optimization
1 Introduction
EAs are a family of population-based, stochastic adaptation models that simulate mechanisms and patterns observed in evolutionary biology and the natural ecosystem. The central idea in these algorithms is to allow a population of artificial creatures—e.g., strings that represent candidate solutions to a problem or programs that perform a given task—to build themselves up through an artificial evolutionary process that aims to approximate what mother-nature does. Over the years, various types of EAs have been proposed, including Evolution Strategies (ES)[1], Evolutionary Programming (EP), Genetic Algorithms (GAs) [2], and Genetic Programming (GP) [3]. The key differences between these algorithms lie in the type of representation used and how self-adaptation is approximated through search operators [4]
By design, EAs are particularly well-suited to solving optimization problems, especially those involving complex, nonlinear, or dynamic environments. This realization can be traced back to the pioneering work of computer scientists in the 1950s and 1960s, who first explored the potential of evolutionary processes as optimization tools for engineering problems [5], and since then, EAs have evolved into a powerful class of computational techniques, capable of addressing diverse problems across science, engineering, and industry.
A system undergoing a natural evolutionary process must continuously compute decisions to carry itself from its current state to a more fit state. The computation of these decisions relies on two complementary sub-processes: exploration and exploitation. Exploration involves the search for resources that may enhance the system chances of survival, and improvement while, exploitation focuses on optimizing the use of known resources to maximize immediate benefits/state-transition. For example, in biological systems, exploration can manifest as genetic mutations, behavioral changes, or the colonization of new habitats, all of which contribute to the genetic diversity and resilience of a population ([6], [7]), while exploitation is essential for refining existing traits or behaviors of a population that have proven successful, thereby enhancing the overall system’s fitness and performance in its current environment.
The interplay between exploration and exploitation is crucial for the long-term success of any evolutionary system. A system that leans too heavily on exploration may become inefficient, wasting resources on unproductive paths, while one that overly focuses on exploitation may become stagnant, failing to adapt to new challenges ([8], [9]). Thus, an effective evolutionary strategy requires a dynamic balance between these two processes to allow the system to adaptively respond to both immediate needs and future uncertainties ([2], [10]).
The same requirement holds for a robust EA. However, unlike nature, which intelligently manages this dilemma, machines do not achieve it inherently. Therefore, human algorithm designers must carefully craft mechanisms that enable EAs to explore the solution space effectively while exploiting promising solutions to ensure convergence, efficiency and optimization.
In the context of algorithm design, the exploitation-exploration dilemma comes down to the following question: How could the algorithm’s capacity for exploration be maximized without compromising its ability to exploit promising solutions? On one hand, there is a desire for the algorithm to thoroughly explore the feasible solution space to ensure comprehensive coverage and the potential discovery of optimal solutions. On the other hand, it is equally crucial to prevent an overamplification of randomness within the algorithm, as this can impede convergence and degrade overall performance.
A review of the literature indicates that the current approach practiced in the literature for addressing this dilemma in EAs primarily centers around fine-tuning and adjusting traditional parameters within EAs. These traditional parameters include: the selection of evolutionary operators, operator probabilities, population size, and others. In this context, researchers often engage in the development of new operators internal to the algorithm and experiment with other various parameter configurations in pursuit of optimizing algorithm performance. And indeed, while specific tuning configurations can lead to improvements, like the selection of particular EA operators outperforming others in specific contexts, the progress achieved through this approach typically exhibits an incremental nature. Furthermore, we argue that this approach inherently faces limitations in delivering effective, intelligible, and noticeable advancements in addressing the dilemma in question.
This paper proposes a novel human-centered two-phase search framework (HCTPS) to resolve this dilemma, and in which, exploration is maximized, but not at the expense of the exploitation potential of the algorithm. The framework consists of two distinct phases: the global search phase, in which the algorithm distributes its internal exploitation-exploration potential across the entire search cube $I$, followed by a local search phase, in which a sequence of selected sub-cubes $\{I_{i}\}$ from the original search cube $I$ (where $i=1,2,\dots,n$) undergoes further intensive sequential exploration. The rigorous implementation of the HCTPS is supported by the introduction of the human-centered search space control parameter (HSSCP), which allows the decision-maker to dynamically specify and adjust the sequence of sub-cubes to explore in the local search phase.
To demonstrate the effectiveness of HCTPS, we integrate the it with the canonical GA and evaluate its performance on 14 well-known benchmark unconstrained optimization problems. The results demonstrate the framework’s ability to enhance exploration while preserving the GA’s inherent robust exploitation, offering a scalable and generalizable solution for various search algorithms.
The remainder of this paper is designed to be as self-contained as possible, with an emphasize on clarity and the use of analogies for intuitive and comprehensive explanation: the next section presents a comprehensive overview of the canonical GA, including its key operators, terminology, and parameters. Section (3) illustrates the inherent limitations of traditional approach in managing the exploration-exploitation trade-off. Section (4) introduces the proposed HCTPS framework, detailing its conceptual design, the two-phase search process, and the role of the human-centered search space control parameter (HSSCP) in enabling dynamic and adaptive exploration, supported by a relatable analogy for intuitive accessibility. Section (5) evaluates the effectiveness of the HCTPS framework through experimental results obtained from benchmark unconstrained optimization problems, comparing its performance with the traditional approach, and ends with a reflective discussion, while section (6) concludes the paper and outlines potential directions for future research.
2 The Canonical Genetic Algorithm (GA)
GAs are among the most important and widely applied EAs for solving complex computational tasks ([11], [12]). The first variant of GAs was introduced by Holland in the 1960s and later refined in the 1970s through collaborations at the University of Michigan. As the name suggests, GAs draw inspiration from Darwinian principles of evolution, particularly natural selection, genetic variation, and the survival-of-the-fittest paradigm ([13]).
In this work, we adopt the function-optimizer interpretation of the GA and specifically implement Holland’s original formulation, commonly referred to in the literature as the simple or canonical GA. The progression of the GA is organized into six key steps, implemented through specific GA operators that collectively shape its evolutionary cycle. These steps are detailed below.
2.1 Terminology, Key Steps, and Operators
Step 0. Encoding
The algorithm begins by generating a set of potential solutions, $S_{r}$, to the problem at hand of cardinality $N$, chosen uniformly at random from within the feasible solution set $S$. Each solution in $S_{r}$ is then encoded (or represented)  [^1] Throughout this paper, we use the terms encoding and representation interchangeably into a finite string of bits of fixed length $L$  [^2] The length of a string is the number of its constituent bits..
Formally, the encoding process is described by the bijective map:
ENCODE: $S_{r}\to X$,
Where $X$ is the set of all possible strings with length $L$ in the chosen representation system. For example, with binary encoding of length $L$, $X$ is the set of all binary strings of length $L$, and the cardinality of $X$ is $|X|=2^{L}$. The selection of the encoding system—whether binary, ternary, decimal, alphabetic, or another format— depends on the characteristics of the problem and must be selected carefully to ensure semantic accuracy and efficiency in representing the solutions.
Remark 1.
Alternatively, one may begin with a randomly generated subset of strings of length $L$ from a chosen representation system (e.g., binary, ternary, etc.) and then apply a DECODE function to map these strings back into the original feasible set $S$.
Terminology:
  • Each encoded string is referred to as chromosome.
  • Each bit constituting the chromosome is called a gene and a population is a collection of chromosomes.
  • We call the set $X$ the representation set, while the base elements of the representation system (e.g. binary {0,1}, ternary {0,1,2}, ..etc) are called representative elements.
Step 1. Evaluation
Once the population has been encoded, the fitness of each chromosome is evaluated using a non-negative function derived from the main objective function to be optimized called the "fitness function". The fitness function assigns a numerical fitness value to each chromosome, serving as a quantitative measure of its performance and quality within the current population. The idea behind assigning fitness values is to discriminate good from bad quality solutions.
Step 2. Selection
Based on the fitness values, a probability measure is derived to determine the likelihood of each potential solution (chromosome) being selected as a parent for reproduction in the next step. The aim here is to reward fitter chromosomes with a survival advantage, that is, a higher probability of contributing their genetic material to the next generation, and to penalize the less fit, making them more likely to die out/get excluded in the process. This aligns with the ”survival of the fittest” principle.
In nature, however, unfit genes are not always immediately eliminated. Some may become recessive, persisting within the population for extended periods. This persistence allows for the possibility that future environmental changes or mutations may transform these genes into beneficial traits. As such, there is a trade-off between promoting fitter chromosomes with advantageous traits and preserving genetic diversity within the population.
GA designers acknowledge this trade-off by employing selection operators that balance the need for promoting fitness with maintaining population diversity. These operators are typically classified based on how they assign probabilities of selection—commonly referred to as selection pressure—to chromosomes within the population. Selection methods can be broadly divided into two categories: roulette-wheel selection and ordinal methods.
Step 3. Recombination
In the recombination step, the GA combines bits and pieces of two or more parental chromosomes through the crossover operator to generate new hopefully better offspring chromosomes with a pre-defined probability $p_{c}$  [^3] Further details can be found in the next section. This operator aims to create offspring that inherit desirable traits from their parents while introducing novel combinations of genetic material.
The specific definition and implementation of the crossover operator can vary, and its effectiveness greatly depends on designing a suitable recombination mechanism. Various approaches exist, such as single-point crossover, multi-point crossover, uniform crossover, and other specialized methods tailored to the problem domain.
The fundamental principle behind recombination is to ensure that the offspring are distinct from the parents. This allows the algorithm to explore new regions of the solution space while preserving key features of the parent chromosomes that contribute to fitness.
Step 4. Mutation
Unlike the crossover operator which operates on two or more parental chromosomes, the mutation operator locally but randomly, modifies a chromosome according to a pre-defined probability $p_{m}$ which is often small.
There are many variations of the mutation operators. The common aspect shared by all the variations is the involvement of one or more changes that are made to the genes of the chromosomes. In other words, mutation performs a random walk in the vicinity of a candidate solution, adding a level of stochasticity and diversity to the GA.
Step 5. Replacement
Finally comes the last step, after the recombination and mutation step which results in forming a new population, we need to introduce the new generated population in replacement with the old one through the replacement operator. There are many replacement operators to help achieve this task such as elitist replacement, generation-wise replacement, and steady-state replacement. These operators determine how the chromosomes from the new population are integrated into the next generation, ensuring the continuous evolution of the population over successive generations.
2.2 Parameters
A parameter in the context of an optimization algorithm is a configuration variable that influences the algorithm’s search navigation, behaviour, performance, and convergence properties. Formally, for an algorithm $\mathcal{A}$, operating on a solution space $\mathcal{X}$, a parameter $\theta\in\Theta$, where $\Theta$ is the parameter space, serves as an input to configure the algorithm but is not subject to optimization during its execution. This can be expressed as:
$\mathcal{A}:(\mathcal{X},\theta)\rightarrow\mathcal{Y},$
where $\theta$ adjusts the mapping between inputs $\mathcal{X}$ and outputs $\mathcal{Y}$.
Parameters can be classified based on their impact, update strategies, and temporal behaviour.
  • Temporal behaviour: Some parameters act as internal configuration variables and evolve during the optimization process, while others act as external configuration variables set before the optimization process, and remain static throughout the algorithm’s execution. Examples:
    • Internal configuration variable: The mutation probability $p_{m}$ in a GA can be set internally based on the algorithm’s feedback. If the algorithm is not converging well, the mutation rate can be increased to introduce more exploration in the search space.
    • External configuration variable: The population size is set externally by the practitioner before running the GA and remains constant throughout the optimization process.
  • Updating strategy: Parameters in an optimization process may either self-update based on feedback from the algorithm or be fixed by the practitioner. Examples:
    • Self-updated parameter: In a GA, $p_{m}$ can be dynamically adjusted based on the fitness improvement of the best solutions found. If the fitness improves rapidly, $p_{m}$ can be reduced to focus more on exploiting the promising solutions, or in a self-adaptive penalty-based GA where the penalty factor gets self-updated.
    • Practitioner-updated parameter: The maximum number of iterations, and the selection of the GA operators are typically specified by the practitioner and remains fixed throughout the optimization process.
  • Impact: Some parameters have more impact than others on the algorithm’s output/behavior in the sense that a small change in their value affects the algorithm’s output, even with slight adjustments, while others may require larger changes to produce noticeable effects.
Below, we analyze the traditional GA parameters in regards to their impact on the exploration-exploitation of the search navigation.
  1. 1.
    Population size Population size is a critical parameter for balancing the exploration and exploitation dynamics of a GA. A small population size limits exploration by reducing diversity, which increases the risk of premature convergence to suboptimal solutions ([14]). Conversely, a larger population size promotes exploration by sampling a broader range of the solution space, reducing the likelihood of stagnation in local optima. However, this comes at the cost of slower convergence, which may hinder the algorithm’s exploitation of promising solutions.
  2. 2.
    Crossover probability $p_{c}$. The crossover operator is the primary search mechanism GAs and one of their defining features compared to other EAs ([15]). The crossover probability $p_{c}$ controls the likelihood that two parent solutions will undergo crossover to produce offspring. A higher $p_{c}$ accelerates convergence by intensifying the recombination of genetic material, potentially leading to faster exploitation of promising solutions. However, excessive crossover may reduce population diversity, hindering the algorithm’s ability to explore the solution space effectively.
    Example: If $p_{c}=1$ and the population consists of $N$ chromosomes, every pair of chromosomes will undergo crossover. In contrast, if $p_{c}=0.5$, there is a $50\%$ chance for any given pair to participate in crossover.
  3. 3.
    Mutation probability $p_{m}$. The mutation probability $p_{m}$ dictates the likelihood of a chromosome undergoing random gene-modifications. Higher $p_{m}$ enhance exploration by introducing randomness, which helps the algorithm escape local optima and discover new regions of the solution space. However, excessive mutation can disrupt favorable solutions, destabilizing the algorithm and reducing convergence efficiency. Conversely, lower $p_{m}$ promotes exploitation by preserving high-quality solutions but may decrease diversity, increasing the risk of premature convergence.
    Example: If $p_{m}=0.1$, there is a $10\%$ chance for each chromosome in the population to undergo mutation. In a population of $N$ chromosomes, approximately $\frac{N}{10}$ of chromosomes are expected to be selected for mutation. However, due to the stochastic nature of the process, the actual number of mutations may vary.
  4. 4.
    Selection of genetic operators. The choice of GA operatore—including selection, crossover, mutation, and replacement—determines the descriptive content (or the logic) of the mechanism by which individuals are chosen for reproduction, how genetic material is recombined, and how it is altered to produce new solutions.
    Example: Consider optimizing a function where the objective is to minimize a complex, multimodal surface. Descriptively different selection operators will have different impact on the exploitation-exploration dynamics.
    • Elitism selection: In each generation, the best-performing solutions (e.g., top 10%) are guaranteed to survive into the next generation. This promotes exploitation by preserving high-quality solutions but risks reducing diversity if overused, potentially causing premature convergence to local optima.
    • Tournament selection: Groups of solutions compete, and the best solution in each group is selected for reproduction. This balances exploration and exploitation by favoring fitter solutions while still allowing less fit ones a chance to propagate, maintaining diversity.
    • Random selection: Individuals are chosen without regard to fitness. This maximizes exploration, ensuring broad sampling of the solution space, but sacrifices exploitation, making it unlikely to focus on improving promising solutions.
  5. 5.
    Representation. The representation of solutions determines how the genetic information is encoded and decoded for the problem at hand. The choice of representation directly affects the search space, the feasibility of generated solutions, and the algorithm’s exploration and exploitation dynamics. A well-suited representation ensures efficient exploration and accurate mapping of candidate solutions.
    Example: The Traveling Salesman Problem (TSP) requires finding the shortest route that visits a set of cities and returns to the starting city. Two common representations are:
    • Binary Representation: Here each gene represents whether a city is visited or not. For example, a chromosome might be represented as [1, 0, 1, 1, 0] indicating that the first, third, and fourth cities are visited. The order of the cities is not explicitly represented in this representation.
    • Permutation Representation: Here, we choose a permutation representation where each gene represents the order in which the cities are visited. For example, a chromosome might be represented as [3, 1, 4, 2, 5] indicating that the cities are visited in the order of 3, 1, 4, 2, and 5.
    Impact on the exploration-exploitation dynamics:
    • The choice of binary representation limits the search space to a subset of all possible permutations which negatively affects the algorithm’s exploration. Since the order of the cities is not explicitly represented, the GA needs to incorporate additional mechanisms to generate valid tours. This representation may require more complex decoding mechanisms and genetic operators to ensure the generated solutions are feasible tours. However, it can be advantageous in reducing the search space complexity and improving computational efficiency, especially for large TSP instances.
    • The choice of permutation representation allows for a direct mapping of the chromosome to a valid tour. The search space includes all possible permutations of the cities, which is a larger search space compared to the binary representation. However, the permutation representation simplifies the decoding process and ensures that all generated solutions are valid tours. This representation may require more computational resources due to the larger search space, but it allows for a more straightforward exploration of different tour configurations.
  6. 6.
    Number of fitness function evaluations. Increasing the number of evaluations allows the algorithm to gather more information about the problem space, enhancing its ability to explore diverse regions and identify potential solutions in a meaningful way. At the same time, this provides opportunities to refine high-performing solutions, improving accuracy and solution quality through exploitation. However, excessive evaluations can significantly increase computational costs, slowing down the algorithm, while too few evaluations may result in premature convergence and suboptimal solutions.
    An analogy can be drawn to a traveler lost in a dark forest with a limited supply of oil for their lamp. Adding more oil brightens the lamp, allowing for better and meaningful navigation and exploration of the forest. However, the faster the oil is used, the less is left for future needs. Similarly, increasing fitness evaluations enhances both exploration and exploitation quality but requires careful balancing to optimize the use of computational resources. Striking this balance is essential to ensure the algorithm is efficient while achieving high-quality results.
2.3 Some Key Insights
  1. 1.
    Efficient search navigation in a GA fundamentally relies on achieving a balanced exploration-exploitation dynamic.
  2. 2.
    Exploitation and exploration in a GA can be addressed on two distinct levels: the semantic level and the intensity level. The semantic level focuses on the descriptive content of the mechanisms or strategies used to achieve exploration and exploitation (the "how"). In contrast, the intensity level determines the degree to which these mechanisms are applied (the "how much"). Together, these levels shape the algorithm’s ability to balance searching diverse areas of the solution space with refining promising solutions.
  3. 3.
    The traditional internal operators (selection, crossover, mutation, replacement) and parameters (e.g., population size, crossover probability, mutation rate) collectively define the GA’s potential for exploration and exploitation.
  4. 4.
    Internal exploration and exploitation processes in a GA, driven by traditional operators and parameters, are inherently conflicting. Optimizing one often diminishes the other
  5. 5.
    The structure of GA tells us that the algorithm operates and moves within the representation set $X$. The cardinality of $X$ is $|X|=K^{L}$ where $K$ is the number of representative elements and $L$ is the length of the string.
  6. 6.
    The structure of GA is abstract enough to give the user flexibility in defining their own operators and parameters, enabling customization and adaptation to specific problem domains.
  7. 7.
    Parameter tuning in optimization algorithms is a crucial task, as it significantly impacts algorithm behavior, convergence, and solution quality. It remains an open question in research to continuously discover new parameters tailored to specific problem domains, leading to enhanced optimization outcomes and novel insights.
3 Managing the Exploration-Exploitation Dilemma in EAs: The Current Approach
Building upon the seventh paragraph in the introduction section, this section analyzes the limitations of traditional approach for enhancing exploration without compromising exploitation in GAs. Although GAs are used as a representative case, these limitations are applicable to the broader class of EAs.
To illustrate these limitations, let us consider the idealized case in which no point in the search space is revisited throughout a single execution of the algorithm.
Increasing the exploration potential of a GA requires fulfilling two necessary conditions: (1) increasing the size of the representation set and (2) the size of the population. As mentioned in section (2), the size of the representation set within which the GA operates is $K^{L}$, where $K$ is the base of the representation system (e.g., binary (2), ternary (3), decimal (9), etc.) and $L$ is the length the representative string of each potential solution. So, increasing the size of the representation set, for some fixed representation system, implies increasing the value of $L$.
Assume that the algorithm explores $N$ solutions per iteration, and given the assumption that no point in the search will be visited twice (which is in practice, not guaranteed to happen) this would imply the need for $m$ number of iterations where $m$ satisfies $\left|m-\frac{K^{L}}{N}\right|\leq\epsilon$ and $\epsilon=\min\left\{y\in\mathbb{Z^{+}}\mid\left|m-\frac{K^{L}}{N}\right|\leq y\right\}$ for the algorithm to explore the whole representation set.
For instance, in binary representation-based GAs, the size of the representation set is $2^{L}$. If $L$ is set to be 20 then we have $10485762^{20}=$ 1048576 different binary strings available to be explored. If $N=100$ then, ideally we would need a prohibitively large number of iterations, $m=10486$ iterations, (or slightly smaller) per run- under the assumption that no point will be visited twice -to explore the whole representation set (or a reasonable portion of it) and with that, we may never reach the best solutions.
It can be inferred that with the addressed approach, enhancing the exploration potential not only entails satisfying conditions (1) and (2) but also demands meticulous tuning of internal GA operators (e.g. crossover + mutation) to facilitate the exploration of new search points as much as possible and minimize the risk of getting trapped in a narrow and suboptimal region of the solution space but without risking it behaving like a random algorithm.
Moreover, several challenges arise immediately:
  • The first challenge lies in the significant computational burden incurred when trying to meet conditions (1) and (2). Increasing the size of the representation set and the population requires substantial computational resources, which can limit the efficiency of the algorithm.
  • The second challenge stems from the inherent unintelligibility of the interdependencies between internal GA parameters and operators which further complicates the understanding and prediction of their combined effects on the exploitation-exploration trade-off thereby, making their tuning process difficult.
  • The third challenge is that even if conditions (1) and (2) are met, and a selection of good GA operators was employed, there are no guarantees that the algorithm will reach promising solutions. It is possible that some promising niches may not be included within the bounds of the workspace  [^4] The pre-image of the representation set under the DECODE map, if we are following the DECODE-ENCODE strategy, or the domain of the ENCODE map, if we are following the ENCODE-DECODE strategy of the algorithm, leading to the risk of overlooking optimal solutions.
The limitations identified in the current approach are not exclusive to GAs but are inherent to the broader class of EAs. This is because all EAs share a common foundation: they are population-based methods that encode candidate solutions using structured representation systems and rely on stochastic operators to guide the search process.
Given these challenges, it is clear that an alternative approach is needed to address the exploitation-exploration dynamics more effectively.
4 A Novel Human-Centered Two-phase Search Framework (HCTPS)
In this section, we propose a human-centered framework, designed to enhance the exploration of the search space while preserving the exploitation capability of an EA, ensuring it does not devolve into random behavior. This framework is structured as a two-phase search process, consisting of:
  1. 1.
    Global Search Phase: In this phase, a broad exploration of the entire search cube $I$ is conducted. In this, the algorithm’s internal exploration-exploitation potential is distributed across the whole search space. However, this phase alone may overlook certain promising regions or niches within $I$, necessitating a subsequent localized search to address this limitation.
  2. 2.
    Local Search Phase: In this phase, a sequence of selected sub-cubes $\{I_{i}\}$ from the original search cube $I$ (where $i=1,2,\dots,n$) undergoes further intensive sequential exploration. In this, the algorithm focuses its internal exploration-exploitation potential over each sub-cube of the generated sequence.
    The rigorous implementation of the local search phase is achieved through the introduction and utilization of the human-centered search space size control parameter (HSSCP) which serves as a defining mechanism for the selection and exploration of the $n$ specific sub-cubes within $I$.
4.1 The Human-Centered Search Space Size Control Parameter (HSSCP)
The HSSCP is an external configuration variable to an EA that represents a selection of a sequence of sub-cubes $\{I\}_{i=1}^{n}\subset I$ for targeted exploration during the local search phase.
The following steps outline how the decision-maker sets, tunes, and updates the HSSCP:
  1. 1.
    Rule Specification: Before the local search phase begins, the decision-maker establishes a rule to generate the sequence of sub-cubes $\{I\}_{i=1}^{n}$.
  2. 2.
    Sequence Generation: Subsequently, the decision-maker generates the first $n$ terms of this sequence according to the specified rule, where each term corresponds to a localized region of interest that the algorithm will intensively explore.
Throughout this process, the decision-maker retains complete flexibility in determining the number of sub-cubes $n$ to be explored. They also have the autonomy to generate additional terms in the sequence as needed, enabling real-time adjustments based on the optimization problem’s complexity, insights gained during exploration, and individual preferences.
4.2 HCTPS: Pseudo-Algorithm
The presented pseudo-algorithm outlines the general framework of the two-phase search process with HSSCP integration.
The specific implementation details may vary based on the algorithm $g$ used, and the characteristics of the output data, including whether it constitutes a single solution set, a multi-solution set (such as a Pareto set in case of multiobjective optimization), or descriptive statistics measures .. etc.
Algorithm 1 HCTPS Algorithm
Input
     Original search cube $I$, HSSCP sequence length $n$, evolutionary algorithm $g$,
     stopping criteria: decision-maker satisfaction $T_{\text{satisfactory}}$
Output
     The best results found during the entire exploration process to $T_{\text{satisfactory}}$
Step 1: Global Search Phase:
 a) Run $g$ across the entire search cube $I$ to perform a broad exploration experiment.
 b) Record the output results.
Step 2: Local Search Phase with HSSCP:
 a) Specify the rule to generate a sequence of sub-cubes $\{I\}_{i=1}^{n}$ within $I$,
 and generate the first $n$ terms of the sequence, where each term represents a localized region of interest
 b) Run $g$ through each sub-cube to conduct an intensive exploration.
 c) Record the output results of each $i$ exploration experiment.
Step 3: Additional Sub-Cubes Generation:
while the stopping criteria $T_{\text{satisfactory}}$ is not met do
      a) Generate $n_{\text{additional}}$ new terms in the sequence and go to step 2c.
end while
Remark 2.
The value of $n_{\text{additional}}$ represents the number of additional terms in the sequence that can be generated after the $n$ exploration experiments are carried out in real-time. This allows the decision-maker to dynamically adjust the exploration based on insights gained, problem complexities, and satisfaction level.
4.3 Visualizing the HCTPS with a Helicopter Analogy
Imagine you are a photographer and you want to capture a breathtaking landscape with a camera equipped with a fixed zoom capability. You take a helicopter to get a bird’s-eye view of the entire scenery from a height of 3000 feet. You admire the vastness of the scenery but realize that crucial details are missing from this altitude. Thereby, you decide to zoom in to be able to explore specific portions of the landscape. Lowering the helicopter, you hover closer to the ground, focusing your camera on different regions to capture the intricacies of each area. As you lower the helicopter and hover closer to the ground while changing your position across the x-y plane, adjusting your camera’s focus on different regions, you will be able to capture new details that were previously hidden from your previous vantage point.
By moving your helicopter strategically to various positions in the x-y-z plane, your camera will be able to capture diverse perspectives of the landscape. Each new position offers a unique view, revealing different aspects of the scenery that were previously hidden from other vantage points, allowing you to compose a more faithful representation of the scenery, compared to the one and only 3000 feet view photoshoot.
This dynamic approach intelligently compensates for the camera’s fixed zoom capability through the human-centered adjustment of the helicopter’s position, allowing exploration of the landscape from various positions and angles. In this scenario, the camera’s fixed zoom capability is analogous to the algorithm’s internal exploitation-exploration fixed potential. The 3000 feet view is analogous to the initial search phase (the global phase) in our proposed approach and this is where traditional EAs stop at thereby, only able to explore the search space from a fixed perspective.
The human-centered strategic position-changing of the helicopter is analogous to the role of the decision-maker in directing the local search phase in 1 where the active participation of the photographer in the strategic adjustment of the helicopter’s position in the x-y-z plane is akin to the active participation of decision-maker in adjusting and tuning the SSCP by generating the sequence of sub-cubes to be explored in the local search phase according to some rule.
5 GA-integrated HCTPS: Experimental Results
To evaluate the effectiveness of the proposed framework, we compare the performance of the standalone GA with the GA integrated into the HCTPS framework described by algorithm (1) where $g$ is the canonical GA on fourteen challenging black-box, unconstrained optimization problems from ([16], [17], [18], [19]), focusing exclusively on their 30-dimensional versions.
These problems exhibit varying levels of difficulty and topological complexity. For instance, functions like F1 (Bent Cigar) are relatively simple and serve as a baseline for performance comparison, while others, such as F10 (Rastrigin) and F11 (Griewank), present significant challenges due to their multimodal nature and the presence of numerous local optima. Moreover, the set contains functions that are specifically designed to be computationally expensive such as F14 (Ackley), thereby simulating real-world scenarios where evaluations of the objective function are costly ([20]), and providing a realistic framework for assessing the effectiveness of various exploration and exploitation EA strategies to navigate complex landscapes ([21]. Detailed discussions of these functions’ properties are provided in ([22], [16]).
Our GA operates with a maximum limit of exact objective function evaluations per run, defined as $\textit{feval}=50\times n$, where $n$ is the dimensionality of the problem. For the 30-dimensional case, $\textit{feval}=1500$. The search cube for all problems is $I=[-100,100]^{30}$, which includes the global minimum for each test problem as well. The GA employs binary encoding, two-point crossover, and tournament selection. Statistical results are derived from 20 independent runs.
All experiments in this research were conducted using MATLAB 2022a (The MathWorks, Inc.) on a personal computer running Microsoft Windows 11, equipped with an 11th Gen Intel(R) Core(TM) i7-1165G7 processor @ 2.80 GHz and 16 GB of RAM. Table (1) presents the results of the GA integrated into the HCTPS framework (HCTPS-GA), while Tables (3, 3, 8, 8) provides a side-by-side comparison between the standalone GA and HCTPS-GA.
ID Function Name Mean Best Worst Median St. Dev Execution Time of 20 Runs (s)
F1 Bent Cigar 4.27603E-38 2.37593E-38 5.79992E-38 4.23246E-38 8.7447E-39 1.252549
F2 Discus 1.16259E-16 4.2676E-20 5.50934E-16 5.60015E-17 1.67102E-16 1.132544
F3 Weierstrass 6.3873E-05 5.45428E-05 7.25862E-05 6.29509E-05 5.08126E-06 8.136591
F4 Modified Schwefel 0.008579635 0.006739378 0.011169685 0.008389775 0.001468878 1.256719
F5 Katsuura 1.13712E-10 9.5848E-11 1.26901E-10 1.14559E-10 7.79472E-12 5.830386
F6 HappyCat 2.840347319 2.840347319 2.840347319 2.840347319 4.07358E-12 1.16361
F7 HGBat 0.500000001 0.500000001 0.500000001 0.500000001 7.64014E-11 1.213726
F8 Expanded Griewank plus Rosenbrock 0.459947694 0.459947694 0.459947694 0.459947694 2.64002E-11 1.429998
F9 Expanded Scaffer’s F6 4.43632E-11 6.56031E-12 9.47957E-11 3.7284E-11 3.07707E-11 1.326823
F10 Rosenbrock’s 2 2 2 2 3.27683E-11 1.200114
F11 Griewank’s 0 0 0 0 0 2.515366
F12 Rastrigin’s 0 0 0 0 0 1.292805
F13 High Conditioned Elliptic 5.42E-20 3.99E-20 7.49E-20 5.54E-20 9.02E-21 1.20E+00
F14 Ackley 1.68E-10 1.47E-10 1.84E-10 1.68E-10 1.18E-11 1.343571
Table 1: Statistical Results of HCTPS-GA
ID HCTPS-GA GA
F1
Mean
Best
Worst
Median
St. Dev
4.27603E-38
2.37593E-38
5.79992E-38
4.23246E-38
8.7447E-39
7.0816E+10
5.7076E+10
9.1041E+10
7.0687E+10
9.5167E+9
F2
Mean
Best
Worst
Median
St. Dev
1.16259E-16
4.2676E-20
5.50934E-16
5.60015E-17
1.67102E-16
1.3119E+8
1.0460E+7
8.4294E+8
1.0484E+7
2.2834E+8
Table 2: Performance Comparison (F1-F2)
ID HCTPS-GA GA
F3
Mean
Best
Worst
Median
St. Dev
6.3873E-05
5.45428E-05
7.25862E-05
6.29509E-05
5.08126E-06
6.3440E-5
5.6442E-5
7.0481E-5
6.2658E-5
3.9426E-6
F4
Mean
Best
Worst
Median
St. Dev
0.008579635
0.006739378
0.011169685
0.008389775
0.001468878
5.9077E+3
4.7541E+3
7.5088E+3
5.8122E+3
6.6643E+2
Table 3: Performance Comparison (F3-F4)
ID HCTPS-GA GA
F5
Mean
Best
Worst
Median
St. Dev
1.13712E-10
9.5848E-11
1.26901E-10
1.14559E-10
7.79472E-12
1.1122E-10
8.9499E-11
1.2968E-10
1.1166E-10
1.1879E-11
F6
Mean
Best
Worst
Median
St. Dev
2.840347319
2.840347319
2.840347319
2.840347319
4.07358E-12
5.1220E+0
2.8251E+0
3.2955E+1
2.8251E+0
6.8450E+0
Table 4: Performance Comparison (F5-F6)
ID HCTPS-GA GA
F7
Mean
Best
Worst
Median
St. Dev
0.500000001
0.500000001
0.500000001
0.500000001
7.64014E-11
6.9444E+4
5.4066E+4
8.4695E+4
7.0732E+4
1.0067E+4
F8
Mean
Best
Worst
Median
St. Dev
0.459947694
0.459947694
0.459947694
0.459947694
2.64002E-11
3.5113E+12
6.6581E+3
2.7423E+13
4.6568E+10
8.1503E+12
Table 5: Performance Comparison (F7-F8)
ID HCTPS-GA GA
F9
Mean
Best
Worst
Median
St. Dev
4.43632E-11
6.56031E-12
9.47957E-11
3.7284E-11
3.07707E-11
3.3533E-1
-4.1880E-1
4.9625E-1
4.6182E-1
2.5453E-1
F10
Mean
Best
Worst
Median
St. Dev
2
2
2
2
3.27683E-11
1.2824E+8
4.1754E+3
6.4375E+8
6.8138E+7
1.6987E+8
Table 6: Performance Comparison (F9-F10)
ID HCTPS-GA GA
F11
Mean
Best
Worst
Median
St. Dev
0
0
0
0
0
1.7259E+1
1.0755E+1
2.2015E+1
1.7394E+1
2.7755E+0
F12
Mean
Best
Worst
Median
St. Dev
0
0
0
0
0
7.0787E+4
4.6881E+4
9.2978E+4
6.9242E+4
1.0849E+4
Table 7: Performance Comparison (F11-F12)
ID HCTPS-GA GA
F13
Mean
Best
Worst
Median
St. Dev
5.42E-20
3.99E-20
7.49E-20
5.54E-20
9.02E-21
6.8974E+4
5.8835E+4
7.9979E+4
6.8824E+4
6.4668E+3
F14
Mean
Best
Worst
Median
St. Dev
1.68E-10
1.47E-10
1.84E-10
1.68E-10
1.18E-11
2.1332E+1
2.1121E+1
2.1544E+1
2.1363E+1
1.0935E-1
Table 8: Performance Comparison (F13-F14)
5.1 Discussion
Throughout these experiments,various configurations of the HSSCP were employed during the local search phase, with each configuration corresponding to a distinct sequence-generative rule. For clarity, details of experiments involving sequence-generative rules that yielded unsatisfactory results in comparison to the global search phase are omitted. Instead, this section focuses on configurations that proved effective for each problem instance. The reason is, unsuccessful attempts in heuristics cannot provide us with information according to which one can infer any result general enough about the function in question or sufficient to theorization. The results of the global search phase are presented in table (1).
For the local search phase, the HSSCP configuration involved the following steps:
  1. 1.
    Lower-dimensional subdivision (dim=3): The cube $I=[-100,100]^{3}$, was subdivided into $82^{3}=$ 8 equal-volume subcubes by splitting each dimension at its midpoint. The resulting subcubes were arranged into a sequence based on a clockwise traversal of the lower layer, followed by the upper layer of $I$ Starting at one corner of the lower layer, the sequence progresses as:
    $\{I_{n}\}_{n=1}^{8}=\begin{aligned} &\big{\{}[-100,0]\times[-100,0]\times[-100,0],[-100,0]\times[-100,0]\times[0,100],\\ &[-100,0]\times[0,100]\times[-100,0],[-100,0]\times[0,100]\times[0,100],\\ &[0,100]\times[-100,0]\times[-100,0],[0,100]\times[-100,0]\times[0,100],\\ &[0,100]\times[0,100]\times[-100,0],[0,100]\times[0,100]\times[0,100]\big{\}}.\end{aligned}$
    Local search was performed within each subcube to identify the region with the best results. For example, for F1, the subcube:
    $\mathcal{S}_{3}=[0,100]\times[-100,0]\times[0,100],$
    was selected
  2. 2.
    Extension to higher dimensions: The selected subcube was extended to 30 dimensions by cyclically replicating its bounds across all dimensions. For FI, for any dimension $x_{i}$, the bounds were:
    $x_{i}\in\begin{cases}[0,100],&\text{if }(i-1)\bmod 3=0,\\ [-100,0],&\text{if }(i-1)\bmod 3=1,\\ [0,100],&\text{if }(i-1)\bmod 3=2.\end{cases}$
    This resulted in the 30-dimensional search subcube:
    $\mathcal{S}_{30}=\prod_{k=1}^{10}\mathcal{S}_{3},$
    where $\mathcal{S}_{3}$ was replicated 10 times to cover all dimensions.
  3. 3.
    Scaling: For most problems, the 30-dimensional subcube was scaled by a factor of $(1/2)^{m}$, where $m>0$ is an integer. For F1, we used $m=80$, resulting in the final search space:
    $\mathcal{S}_{30,\text{scaled}}=\left(\frac{1}{2}\right)^{80}\times\mathcal{S}_{30}.$
The following tables(9, 10) present the selected subcube similar to $\mathcal{S}_{3}$ to F1, for each test function, and the scaling factors used if any respectively.
Table 9: Selected Subcubes for Each Test Function
ID Selected Subcube ($\mathcal{S}_{3}$)
F1 $[0,100]\times[-100,0]\times[0,100]$
F2 $[0,100]\times[-100,0]\times[-100,0]$
F3 $[0,100]\times[0,100]\times[0,1000]$
F4 $[-100,0]\times[0,100]\times[0,100]$
F5 $[0,100]\times[0,100]\times[-100,0]$
F6 $[0,100]\times[-100,0]\times[0,100]$
F7 $[-100,0]\times[-100,0]\times[-100,0]$
F8 $[0,100]\times[0,100]\times[-100,0]$
F9 $[-100,0]\times[0,100]\times[-100,0]$
F10 $[-100,0]\times[0,100]\times[0,100]$
F11 $[-100,0]\times[-100,0]\times[-100,0]$
F12 $[-100,0]\times[0,100]\times[-100,0]$
F13 $[0,100]\times[-100,0]\times[0,100]$
F14 $[0,100]\times[-100,0]\times[-100,0]$
Table 10: Scaling Factors Applied in the Local Search Phase
ID Scaling Factor ($(1/2)^{m}$)
F1 $(1/2)^{80}$
F2 $(1/2)^{40}$
F3 None
F4 $(1/2)^{10}$
F5 None
F6 $(1/2)^{20}$
F7 $(1/2)^{40}$
F8 $(1/2)^{40}$
F9 $(1/2)^{40}$
F10 $(1/2)^{40}$
F11 $(1/2)^{40}$
F12 $(1/2)^{40}$
F13 $(1/2)^{40}$
F14 $(1/2)^{40}$
As demonstrated in Tables (1, 3, 3, 8, 8, 8, 8, 8), the integration of the GA within the HCTPS framework consistently outperformed the standalone GA in terms of best objective values and statistical measures. Through the HCTPS-GA, we managed to find solutions that are within a tight vicinity of the known global optimum for each objective function, and even achieved the exact global optimum for F11 and F12.
Assessment of the HCTPS framework
The HCTPS is evaluated against the traditional approach based on the following key criteria: achieving the goal of maximizing exploration without minimizing exploitation,and the approach of achievement, and the invested cost vs produced value.
  1. 1.
    Goal achievement, and approach: The HCTPS framework effectively resolves the inherent conflict between exploration and exploitation that is characteristic of the traditional framework. This resolution allows for maximized exploration without compromising exploitation which is fundamental to the HCTPS.
  2. 2.
    Cost-benefit analysis: The HCTPS focuses tuning efforts towards the HSSCP, which is the core mechanism of this framework. While the tuning process may require trial and error to identify promising subcubes in the search space, the benefits significantly outweigh the tuning efforts cost. Not only that through the HSSCP, the limitations of the traditional approach listed in section (3), are effectively circumvented, it is the only reliable way to circumvent them. Also it is a more informed decision to invest more tuning efforts in configuring the HSSCP rather than internal EA parameters given the unintelligible interdependencies between internal parameters of the algorithm and the ambiguity of their specific roles in driving exploration versus exploitation internally. Even when specific HSSCP configurations fail to yield satisfactory results to the user, they still provide valuable insights into the search space by identifying regions that may not need further exploration. Furthermore, the intuitive and straightforward nature of fine-tuning the HSSCP enhances its practicality. This makes the HCTPS framework both an effective and efficient alternative to the traditional one for achieving the desired balance between exploration and exploitation.
6 Conclusions
The HCTPS framework represents a significant advancement in addressing the limitations of the traditional approach to managing the exploitation-exploration dynamics in EAs. By resolving the inherent conflict between these two goals, the framework is thus able to maximize exploration without diminishing exploitation, a necessary condition for effective search navigation.
While the fine-tuning of the HSSCP may involve an element of trial and error, this process is neither unique to HCTPS nor a hindrance to its adoption. EAs universally require trial and error experimentation to tailor parameter values to specific problem contexts. The HSSCP, however, offers much more valuable returns regarding maximizing exploration without introducing unnecessary randomization to the structure of the algorithm, in addition to its adaptability and intuitive implementation. Moreover, the ability to parallelize sub-cube testing provides a practical avenue for alleviating the computational burden, enabling concurrent exploration of multiple regions and expediting the search process.
Furthermore, the HCTPS framework applicability is not limited to EAs; it can be applied to any search algorithm as balancing exploration and exploitation is the fundamental principle to effective search.
7 Future Work
Moving forward, promising avenues for research could potentially lie in advancing human-centered tuning techniques within the HCTPS framework. Specifically, efforts can focus on developing generalizable rules for partitioning the search space to enhance exploration potential, particularly in high-dimensional contexts.
Acknowledgments
I sincerely thank the organizers of the SIAM-MDS24 meeting for providing me with the opportunity to present and discuss a poster on my research, which later evolved into this paper.